# Quantum Geometry Induced Kekulé Superconductivity in Haldane phases

### v1

@@ Abstract
Chiral two-dimensional electron gases, which capture the electronic properties of graphene and rhombohedral graphene systems, exhibit singular momentum-space vortices and are susceptible to interaction-induced topological Haldane phases. Here, we investigate pairing interactions in these inversion-symmetric Haldane phases of chiral two-dimensional electron gases. We demonstrate that the nontrivial band topology of the Haldane phases enhances intra-valley (${\bf Q}=\pm 2{\bf K}_{D}$) pair susceptibility relative to inter-valley (${\bf Q}=0$) pair susceptibility, favoring the emergence of a lattice-scale pair-density wave order. When longitudinal acoustic phonons mediate the pairing interaction, the system supports a chiral Kekulè superconducting order. Our findings are relevant to superconductivity in rhombohedral graphene and Kagome metals.
Introduction.—While significant progress has been made in understanding unconventional superconductivity [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] in spin Chern bands of twisted two-dimensional (2D) crystals, the presence of a moiré superlattice often complicates the identification of the underlying mechanisms [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. For instance, it remains unresolved whether the observed superconductivity originates predominantly from band-projected interactions or from novel pairing instabilities intrinsic to topological bands [30, 31, 32, 33, 34]. Alternatively, multilayer rhombohedral graphene (RG) has emerged as a promising platform, displaying superconducting phenomenology [35, 36, 37, 38, 39, 40, 41, 42, 43] that appears to be as rich as that of twisted 2D crystals [44, 45, 46, 47], but without the complexity of a moirè potential. Notably, in RG, superconductivity can even be hosted by a spin-polarized valley Chern band [45, 35, 36, 37, 38, 39, 40, 41, 42, 43], which underscores the necessity to understand the interplay between band geometry and pairing interactions [48].
@@
This enhancement arises from the coherence factors of the Bloch wavefunctions in the Haldane phase [54], which can be visualized using the pseudo-spinor field defined by the Hamiltonian on the Bloch sphere, as illustrated in Fig. 1. In the case of intra-valley pairing as depicted Fig. 1 a), the pseudo-spinors point in the same direction, resulting in constructive interference. In contrast, for inter-valley pairing shown in Fig. 1 b), the pseudo-spinors point in opposite directions, leading to destructive interference and a corresponding suppression of the inter-valley pair susceptibility. When longitudinal acoustic phonons mediate the pairing, the resulting superconducting order acquires a chiral character. The chiral Kekulé superconductivity arises directly from the projected interactions within the topological bands, reflecting the underlying geometric features of the Haldane phase. We discuss the properties of the chiral Kekulé superconductors and the relevance of our findings to multilayer RG.
@@
where the Pauli matrices $\hat{\sigma}_{i}$ act on the pseudospinor $\psi^{\dagger}=(\psi^{\dagger}_{A},\psi^{\dagger}_{B})$, $(A,B)$ are sublattice degrees of freedom to be specified below, $\tau_{z}=\pm 1$ denote the valleys associated with the Dirac points located at the corners of the hexagonal Brillouin zone (BZ) $\pm{\bf K}_{D}$, ${\bf p}=(p_{x},p_{y})$ is the momentum measured from the Dirac points with $\varphi_{{\bf p}}=\tan^{-1}(p_{y}/p_{x})$, and $p=\sqrt{p_{x}^{2}+p_{y}^{2}}$. We choose the $M$-point of the BZ as the origin, giving $\pm{\bf K}_{D}=(\pm 2\pi/3,0)$ and the valley separation ${\bf Q}=2{\bf K}_{D}=(4\pi/3,0)$. Eq. 1 with $m=0$ describes the minimal model of $J$-layer RG [55], with the pseudospinor defined on the top (1,A)(1,\,A) and bottom (J,B)(J,\,B) layers, $\xi_{J}=v^{J}/\gamma_{1}^{J-1}$, and $v=\sqrt{3}/2\gamma_{0}a_{0}\sim 673$ meV nm-1, where $\gamma_{0}\sim 3.16$ eV and $\gamma_{1}\sim 0.445$ eV are the nearest-neighbor intra- and inter-layer hopping parameters [42], respectively, and $a=0.246$ nm is the graphene lattice constant. To highlight the interplay between band geometry and pairing interactions, we neglect trigonal warping for now and address this aspect later.
The Haldane mass term $m\hat{\sigma}_{z}\tau_{z}$ results in a quantum anomalous Hall (QAH) state with a Chern number $J$ [52]. This mass term breaks time-reversal symmetry while preserving inversion symmetry. In contrast, a displacement field produces a sublattice staggered potential $\Delta_{V}\hat{\sigma}_{z}$ breaking the inversion symmetry. The band dispersion $\epsilon_{{\bf p},J,\pm}=\pm(\zeta_{J}^{2}p^{2J}+m^{2})^{1/2}$ near the $\pm{\bf K}_{D}$ points is particle-hole and inversion symmetric. The Bloch wave functions for the conduction band near $\tau_{z}{\bf K}_{D}$ are $|u_{{\bf p},}(\tau_{z}{\bf K}_{D})\rangle=(\cos(\theta_{{\bf p}}/2)e^{-\imath\tau_{z}J\varphi_{{\bf p}}/2},\sin(\theta_{{\bf p}}/2)e^{\imath\tau_{z}J\varphi_{{\bf p}}/2})$ with $\cos(\theta_{{\bf p}})=m/\epsilon_{{\bf p},J,+}$. The singularities of the Bloch states near the Dirac points exhibit frustrated momentum-space pseudospin structure, creating a “fertile ground” for spontaneous symmetry breaking [56, 57, 52].
@@
A series of experiments [58, 32, 33, 59] have provided compelling evidence for interaction-driven Haldane phase in RG systems ranging from bilayer to pentalayer. In particular, the QAH conductance is quantized at $5e^{2}/h$ in pentalayer RG at zero magnetic field [32]. Effective in these states, only one spin contributes to the QAH effect [52]. Our mean-field analysis for spinless C2DEGs (see End Matter) reveals a negative interaction $u_{\perp}<0$ in the $\tau_{\perp}\equiv{\bf\tau}\times\tau_{z}$ channel of the short-ranged valley-dependent interaction leads to a spontaneous QAH state, whereas $u_{\perp}>0$ yields a fully-layer-polarized state. The interaction-induced Haldane mass term $m$ depends on the chirality $J$, which can be significant $m\sim 50$ meV for $J=4,5$ for a hBN encapsulated device with a dielectric constant $\epsilon\sim 5$, for gate screened interactions with gate distance $d=10$ nm. Since mean-field theory tends to overestimate interaction-induced gaps, we take $m\neq 0$ as a tunable parameter.
@@
$\Pi({\bf q})=-\int\frac{d^{2}{\bf k}}{(2\pi)^{2}}\frac{g({\bf k},{\bf k}+{\bf q})}{\xi_{{\bf k}+{\bf q}}+\xi_{-{\bf k}}}\bigg{(}\frac{1+\hat{\bf h}_{{\bf k}+{\bf q}}\cdot\hat{\bf h}_{-{\bf k}}}{2}\bigg{)}\,$, (2)
@@
Away from the Dirac points, the inter-valley coherence factor and corresponding susceptibility become dependent on the ratio of the Haldane gap $m$ to the Fermi energy $\mu$. The intra-valley (S) and inter-valley (D) pair susceptibilities, denoted by $\Pi_{S}$ and $\Pi_{D}$ for the $\sigma_{0}$ channel, can be expressed as
$\Pi_{S}=\frac{-N_{0}}{(1-x^{2})^{\alpha}}\log\bigg{(}\frac{E_{c}}{T}\bigg{)}+\cdots,\quad\Pi_{D}=\frac{1-x^{2}}{2}\Pi_{S},$ (3)
where $\cdots$ denote non-singular terms, $N_{0}=x^{1-\beta}/(2\pi J\zeta_{J}^{\beta}m^{1-\beta})$ defines a characteristic density-of-states scale, $\alpha=1-1/J$, $\beta=2/J$, $x=m/\mu$, $E_{c}\sim\gamma_{1}$ is the high-energy cutoff, and we have set the Boltzmann constant $k_{B}=1$. Therefore, $\Pi_{S}>\Pi_{D}$ throughout the physically relevant range $x\in(0,1]$. This geometric suppression of inter-valley coherence represents a hallmark of pairing in the Haldane Chern bands, distinguishing it from pairing in conventional Fermi liquids.
@@
$\Pi_{S}(\pm{\bf K}_{D})=\frac{-N_{0}}{(1-(x\pm y)^{2})^{\alpha}}\log\bigg{(}\frac{E_{c}}{T}\bigg{)}+\cdots.$ (4)
Meanwhile, the inter-valley pair susceptibility loses its weak-coupling logarithmic divergence due to the breaking of inversion symmetry. Finally, in the limit, y>1xy>1-x and $y<x$, the system remains in the topological phase but hosts only a single Fermi surface at a single valley. This regime corresponds to a “quarter-metal” superconducting phase [42].
@@
where the electron creation operators at ${\bf K}_{D}$ are defined as $c^{\dagger}_{{\bf K}_{D}+{\bf p}}=a^{\dagger}_{{\bf p}}$ and by $c^{\dagger}_{-{\bf K}_{D}+{\bf p},}=b^{\dagger}_{{\bf p}}$ at ${\bf K}_{D}$. Due to inversion symmetry, the interaction matrix elements $G_{i}({\bf p},{\bf p}^{\prime})$, satisfy $G_{1}({\bf p},{\bf p}^{\prime})=2G_{4}({\bf p},{\bf p}^{\prime})$ and $G_{2}({\bf p},{\bf p}^{\prime})=2G_{3}({\bf p},{\bf p}^{\prime})$. Near the Fermi surface $p_{x}=p_{F}\cos(\varphi_{{\bf p}}),p_{y}=p_{F}\sin(\varphi_{{\bf p}})$ and $|{\bf p}|\sim|{\bf p}^{\prime}|\sim p_{F}$, the interaction matrix elements can be expressed as,
@@
$G_{3}$ $=$ $-\frac{G_{0}}{2}\bigg{[}1+\cos(J\Phi)\bigg{]}\bigg{(}1-\frac{m^{2}}{(\xi_{p}+\mu)^{2}}\bigg{)}\,$, (6)
@@
$\dot{g}_{1,l}$ $=$ $\frac{-(1-x^{2})}{2}[g_{1,l}^{2}+g_{2,l}^{2}],\quad\dot{g}_{2,l}=-(1-x^{2})g_{1,l}g_{2,l},$ (7)
$\dot{g}_{4,l}$ $=$ $-g_{4,l}^{2}-g_{3,l}^{2},\quad\dot{g}_{3,l}=-2g_{3,l}g_{4,l},$ (8)
where $g_{i,l}(x)=N_{0}G_{i,l}/(1-x^{2})^{\alpha}$ and $\dot{g_{i,l}}=dg_{i,l}/dt$, where $t=\log(\omega_{D}/T)$ and $\omega_{D}\sim 0.2$ eV denotes the Debye energy. The coherence factor $(1-x^{2})/2$ in Eq. 7, arising from the quantum geometric properties of the Haldane phase, is essential in the suppression of inter-valley pairing. Eq. 7 and Eq. 8 can be solved by making the substitution $g^{intra}_{\pm,l}=g_{4,l}\pm g_{3,l}$, $g^{inter}_{\pm,l}=g_{1,l}\pm g_{2,l}$ satisfying $(g_{+,l}>g_{-,l})$ giving,
$g^{intra}_{\pm,l}=\frac{g_{\pm,l}}{1+g_{\pm,l}t},\quad g^{inter}_{\pm,l}=\frac{g_{\pm,l}}{1+(1-x^{2})g_{\pm,l}t/2},$ (9)
@@
which implies that the critical temperature for the intra-valley (S) pairing, $T^{S}_{c}(x)=\omega_{D}\exp(-1/|g^{intra}_{+,l}|)$, is higher than that of the inter-valley (D) pairing, $T^{D}_{c}(x)=\omega_{D}\exp(-2/((1-x^{2})|g^{inter}_{+,l}|))$. As shown in Fig. 4 b), the ratio $T^{D}_{c}/T^{S}_{c}$ is exponentially suppressed, indicating that the intra-valley pairing dominates over the inter-valley pairing for all values of $x$. This exponential dependence is a direct result of the ratio of the inter- and intra-valley pair susceptibilities $\Pi_{D}/\Pi_{S}$ as derived in Eq. 3.
@@
We can perform the angular momentum decomposition of the interaction matrix elements associated with $G_{4}({\bf p},{\bf p}^{\prime})$ and $G_{3}({\bf p},{\bf p}^{\prime})$ simply by inspection. Including the pair tunnelling term one finds $G_{+,0}=-G_{0}(1-x^{2})$, $G_{+,J}=-G_{0}(1-x)/2$ and $G_{+,-J}=-G_{0}(1+x)/2$. Since $G_{+,-J}>G_{+,0}>G_{+,J}$ for $x_{c}>0.5$ we find that the chiral-$J$ Kekulé superconductor has the largest $T^{S}_{c}$ when $k_{F}\leq(\sqrt{3}m/\xi_{J})^{1/J}$ corresponding to the densities $n_{e}\leq 1/(2\pi)(\sqrt{3}m/\xi_{J})^{2/J}$. Based on the minimal chiral model, we obtain a phase diagram of superconductivity in Fig. 4 a) for chiral 2DEGs in the carrier density and Haldane mass space for different values of $J=3,4,5$ and $6$. To ensure a simply connected Fermi surface, we have assumed that the electron density $n_{e}\geq 0.8\times 10^{12}$ cm-2, sufficiently large [45, 42] to avoid the three-pocket and annulus Fermi surfaces induced by trigonal warping. As detailed in the End Matter, the mean-field critical temperature of the chiral-$J$ Kekulé state is estimated to be in the range $T^{S}_{c}\sim 90.8\textrm{mK}-13.5\textrm{K}$. Due to large superfluid stiffness of C2DEGs in the dispersive regime [62], we anticipate the Berezinskii–Kosterlitz–Thouless (BKT) transition temperature $T^{S}_{BKT}$ to approach the critical mean field temperature $T^{S}_{BKT}\to T^{S}_{c}$.
@@
Figure 4: a) Phase diagram for the C2DEG model of $J$-layer RG with $J=3,4,5$ and 66 in increasing order. The electron density $n_{e}$ is in units of cm-2 and the Haldane mass is in meV. At low densities, the chiral-$J$ Kekulé superconducting channel dominates. b) The ratio of the inter-valley critical temperature ($T_{c}^{D}$) to the intra-valley critical temperature ($T^{S}_{c}$) as a function of $x$. c) Schematic of the Kekulé superconductor on the projected honeycomb lattice, with the superconducting unit cell indicated by the dotted lines.
@@
Interestingly, the presence of a valley degree of freedom permits even-parity pairing channels, including a momentum-independent $s$-wave pairing when $\mu>m$. However, the $s$-wave pairing channel is suppressed as $\mu\to m$ given $G_{+,0}=-G_{0}(1-x^{2})$. This suppression originates from a combination of the Pauli exclusion principle and the quantum geometric properties of the Haldane phase. At the $\pm{\bf K}_{D}$ points, the sublattice and valley degrees of freedom become coincident and act as a single pseudospin, thereby forbidding local (contact) interactions and rendering $s$-wave Kekulé order incompatible. Consequently, this enhances the relative stability of the chiral-$J$ Kekulé order, which emerges as the dominant channel in this regime. Moreover, due to the presence of a non-zero pair tunnelling channel $G_{3,l}$, we expect the chiral-$J$ Kekulé state to be robust to the application of a displacement field, so long as $\Delta_{V}<m$.

### v2

@@ Abstract
Chiral two-dimensional electron gases, which capture the electronic properties of graphene and rhombohedral graphene systems, exhibit singular momentum-space vortices and are susceptible to interaction-induced topological Haldane phases. Here, we investigate pairing interactions in the inversion-symmetric Haldane phases of chiral two-dimensional electron gases. We demonstrate that the nontrivial band topology of the Haldane phases enhances intra-valley (${\bf Q}=\pm 2{\bf K}_{D}$) pair susceptibility relative to inter-valley (${\bf Q}=0$) pair susceptibility, favoring the emergence of a lattice-scale pair-density wave order. When longitudinal acoustic phonons mediate the pairing interaction, the system supports a chiral Kekulè superconducting order at low densities. At higher densities, the phase diagram depends on the parity of the chiral index $J$. For even parity an s-wave Kekulè order appears, while for odd parity we find a ${\bf Q}=0$ valley triplet chiral-$J$ and a valley singlet s-wave order at different densities. Our findings are relevant to superconductivity in rhombohedral graphene and Kagome metals.
While significant progress has been made in understanding unconventional superconductivity [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] in spin Chern bands of twisted two-dimensional (2D) crystals, the presence of a moiré superlattice often complicates the identification of the underlying mechanisms [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. For instance, it remains unresolved whether the observed superconductivity originates predominantly from band-projected interactions or from novel pairing instabilities intrinsic to topological bands [30, 31, 32, 33, 34]. Alternatively, multi-layer rhombohedral graphene (RG) has emerged as a promising platform, displaying superconducting phenomenology [35, 36, 37, 38, 39, 40, 41, 42, 43] that appears to be as rich as that of twisted 2D crystals [44, 45, 46, 47], but without the complexity of a moirè potential. Notably, in RG, superconductivity can even be hosted by a spin-polarized valley Chern band [45, 35, 36, 37, 38, 39, 40, 41, 42, 43, 48], which underscores the necessity to understand the interplay of band geometry and pairing interactions [49].
@@
This enhancement arises from the coherence factors of the Bloch wavefunctions in the Haldane phase [55], which can be visualized using the $\hat{z}-$component of pseudo-spinor field defined by the Hamiltonian on the Bloch sphere, as illustrated in Fig. 1. In the case of intra-valley pairing as depicted in Fig. 1 a), the $\hat{z}-$component pseudo-spinors point in the same direction, resulting in constructive interference. In contrast, for inter-valley pairing shown in Fig. 1 b), the pseudo-spinors point in opposite directions, leading to destructive interference and a corresponding suppression of the inter-valley pair susceptibility. Overall, the pair susceptibility and superconducting phase diagram depend on the parity of the chiral index $J$. When longitudinal acoustic phonons mediate the pairing interaction, the system supports a chiral Kekulè superconducting order for low densities $n_{e}$. For even values of $J$ the system transitions to an s-wave Kekulè superconducting order when $n_{e}>n_{e,2}$, and chiral Kekulè superconducting order appears for $n_{e}>n_{e,2}$. For odd values of $J$, a $\pi$-phase chiral Kekulè superconducting order appears when $n_{e}<n_{e,1}$, while in the density regime $n_{e,1}<n_{e}<n_{e,2}$, we find a ${\bf Q}=0$ valley triplet chiral superconducting state and a valley-singlet s-wave superconductor for $n_{e}>n_{e,2}$. The chiral superconducting order arises from projected interactions within the topological bands, thereby further reflecting the underlying geometric features of the Haldane phase of C2DEGs.
@@
where the Pauli matrices $\hat{\sigma}_{i}$ act on the pseudospinor $\psi^{\dagger}=(\psi^{\dagger}_{A},\psi^{\dagger}_{B})$, $(A,B)$ are sublattice degrees of freedom to be specified below, $\tau_{z}=\pm 1$ denote the valleys associated with the Dirac points located at the corners of the hexagonal Brillouin zone (BZ) $\pm{\bf K}_{D}$, ${\bf p}=(p_{x},p_{y})$ is the momentum measured from the Dirac points with $\varphi_{{\bf p}}=\tan^{-1}(p_{y}/p_{x})$, and $p=\sqrt{p_{x}^{2}+p_{y}^{2}}$. We choose the $M$-point of the BZ as the origin, giving $\pm{\bf K}_{D}=(\pm 2\pi/3,0)$ and the valley separation ${\bf Q}=2{\bf K}_{D}=(4\pi/3,0)$. Eq. 1 with $m=0$ describes the minimal model of $J$-layer RG [56], with the pseudospinor defined on the top (1,A)(1,\,A) and bottom (J,B)(J,\,B) layers, $\zeta_{J}=v^{J}/\gamma_{1}^{J-1}$, and $v=\sqrt{3}/2\gamma_{0}a_{0}\sim 673$ meV nm-1, where $\gamma_{0}\sim 3.16$ eV and $\gamma_{1}\sim 0.445$ eV are the nearest-neighbor intra- and inter-layer hopping parameters [42], respectively, and $a=0.246$ nm is the graphene lattice constant. To highlight the interplay between band geometry and pairing interactions, we neglect trigonal warping for now and address this aspect later.
The Haldane mass term $m\hat{\sigma}_{z}\tau_{z}$ results in a quantum anomalous Hall (QAH) state with a Chern number $J$ [53]. This mass term breaks time-reversal symmetry while preserving inversion symmetry. In contrast, a displacement field produces a sublattice staggered potential $\Delta_{V}\hat{\sigma}_{z}$ breaking the inversion symmetry. The band dispersion $\epsilon_{{\bf p},J,\pm}=\pm(\zeta_{J}^{2}p^{2J}+m^{2})^{1/2}$ near the $\pm{\bf K}_{D}$ points is particle-hole and inversion symmetric. The Bloch wave function for the conduction band near ${\bf K}_{D}$ is $|u_{{\bf p}}({\bf K}_{D})\rangle=(\cos(\theta_{{\bf p}}/2)e^{-\imath\tau_{z}J\varphi_{{\bf p}}/2},\sin(\theta_{{\bf p}}/2)e^{\imath\tau_{z}J\varphi_{{\bf p}}/2})$ with $\cos(\theta_{{\bf p}})=m/\epsilon_{{\bf p},J,+}$, with inversion symmetry giving $|u_{-{\bf p}}(-{\bf K}_{D})\rangle=\sigma_{x}|u_{{\bf p}}({\bf K}_{D})\rangle$. The singularities of the Bloch states near the Dirac points exhibit frustrated momentum-space pseudospin structure, creating a “fertile ground” for spontaneous symmetry breaking [57, 58, 53].
@@
A series of experiments [59, 32, 33, 60] have provided compelling evidence for interaction-driven Haldane phase in RG systems ranging from bilayer to pentalayer. In particular, the QAH conductance is quantized at $5e^{2}/h$ in pentalayer RG at zero magnetic field [32]. Effective in these states, only one spin contributes to the QAH effect [53]. Our mean-field analysis for spinless C2DEGs (see End Matter) reveals a negative interaction $u_{\perp}<0$ in the ${\bm{\tau}}_{\perp}\equiv{\bm{\tau}}\times\tau_{z}$ channel of the short-ranged valley-dependent interaction leads to a spontaneous QAH state, whereas $u_{\perp}>0$ yields a fully-layer-polarized state. The interaction-induced Haldane mass term $m$ depends on the chirality $J$, which can be significant $m\sim 50$ meV for $J=4,5$ for a hBN encapsulated device with a dielectric constant $\epsilon\sim 5$, for gate-screened interactions with gate distance $d=10$ nm. Since mean-field theory tends to overestimate interaction-induced gaps, we take $m\neq 0$ as a tunable parameter.
@@
$\Pi({\bf q})=\int\frac{d^{2}{\bf k}}{(2\pi)^{2}}\frac{g({\bf k},{\bf k}+{\bf q})}{\xi_{{\bf k}+{\bf q}}+\xi_{-{\bf k}}}\bigg(\frac{1+\hat{\bf h}_{{\bf k}+{\bf q}}\cdot\hat{\bf h}_{-{\bf k}}}{2}\bigg),$ (2)
@@
Away from the Dirac points, the intra-valley (S) and inter-valley (D) pair susceptibilities, denoted by $\Pi_{S}=\Pi({\bf Q}=2{\bf K}_{D})$ and $\Pi_{D}=\Pi({\bf Q}=0)$, dependent on the ratio of the Haldane gap $m$ to the Fermi energy $\mu$. They are different for even and odd values of the chiral index $J$, and can be expressed as
$\Pi_{a}=\frac{N_{0}}{(1-x^{2})^{\alpha}}\log\bigg(\frac{E_{c}}{T}\bigg)f_{a}(x)+\cdots,$ (3)
where $\cdots$ denote non-singular terms, $a=S(D)$ denotes the intra- and inter-valley pairing, $N_{0}=x^{1-\beta}/(2\pi J\zeta_{J}^{\beta}m^{1-\beta})$, denotes the characteristic density-of-states, with $\alpha=1-1/J$, $\beta=2/J$, $E_{c}\sim\gamma_{1}$ is the high-energy cutoff, and we have set the Boltzmann constant $k_{B}=1$. The intra-valley (S) and inter-valley (D) pair susceptibilities, depend on $x=m/\mu$ through the functions $f_{S}(x)$ and $f_{D}(x)$. For even chirality, $f_{S}(x)=1$ and $f_{D}(x)=(1-x^{2})/2$, while for odd chirality $f_{S}(x)=x^{2}$ and $f_{D}(x)=(1-x^{2})/2$. Therefore, for even values of $J$, $\Pi_{S}>\Pi_{D}$ throughout the physically relevant range $x\in(0,1]$. Alternatively, for odd values of $J$, $\Pi_{S}>\Pi_{D}$ requires $x>1/\sqrt{3}$. In both cases, the geometric suppression of inter-valley coherence represents a hallmark of pairing in the Haldane Chern bands.
@@
$\Pi_{S}(\pm 2{\bf K}_{D})=\frac{-N_{0}}{(1-(x\pm y)^{2})^{\alpha}}\log\bigg(\frac{E_{c}}{T}\bigg)f_{S}(x\pm y)+\cdots.$ (4)
Meanwhile, the inter-valley pair susceptibility loses its weak-coupling logarithmic divergence due to the breaking of inversion symmetry. Finally, in the limit y>1xy>1-x and $y<x$, the system remains in the topological phase but hosts only a single Fermi surface at a single valley. This regime corresponds to a “quarter-metal” superconducting phase [42, 48].
@@
where the electron creation operators at ${\bf K}_{D}(-{\bf K}_{D})$ are defined as $c^{\dagger}_{{\bf K}_{D}+{\bf p}}=a^{\dagger}_{{\bf p}}$ and $c^{\dagger}_{-{\bf K}_{D}+{\bf p},}=b^{\dagger}_{{\bf p}}$. Due to inversion symmetry, the interaction matrix elements $G_{i}({\bf p},{\bf p}^{\prime})$, satisfy $G_{1}({\bf p},{\bf p}^{\prime})=2G_{4}({\bf p},{\bf p}^{\prime})$ and $G_{2}({\bf p},{\bf p}^{\prime})=2G_{3}({\bf p},{\bf p}^{\prime})$. Near the Fermi surface $p_{x}=p_{F}\cos(\varphi_{{\bf p}}),p_{y}=p_{F}\sin(\varphi_{{\bf p}})$ and $|{\bf p}|\sim|{\bf p}^{\prime}|\sim p_{F}$, the interaction matrix elements can be expressed as,
@@
$G_{3}$ $=$ $-\frac{G_{0}}{2}\bigg[1+(-1)^{J}\cos(J\Phi)\bigg]\bigg(1-\frac{m^{2}}{(\xi_{p}+\mu)^{2}}\bigg),$ (6)
@@
$\dot{g}_{1,l}$ $=$ $-[g_{1,l}^{2}+g_{2,l}^{2}]f_{S}(x),\quad\dot{g}_{2,l}=-2f_{S}(x)g_{1,l}g_{2,l},$ (7)
$\dot{g}_{4,l}$ $=$ $-[g_{4,l}^{2}-g_{3,l}^{2}]f_{D}(x),\quad\dot{g}_{3,l}=-2f_{D}(x)g_{3,l}g_{4,l},$ (8)
where $g_{i,l}(x)=N_{0}G_{i,l}/(1-x^{2})^{\alpha}$ and $\dot{g_{i,l}}=dg_{i,l}/dt$, where $t=\log(\omega_{D}/T)$ and $\omega_{D}\sim 0.2$ eV denotes the Debye energy. Eq. 7 and Eq. 8 can be solved by making the substitution $g^{intra}_{\pm,l}=g_{4,l}\pm g_{3,l}$, $g^{inter}_{\pm,l}=g_{1,l}\pm g_{2,l}$ which gives
$g^{intra}_{\pm,l}=\frac{g^{intra}_{\pm,l}}{1+f_{S}(x)g^{intra}_{\pm,l}t},\quad g^{inter}_{\pm,l}=\frac{g^{inter}_{\pm,l}}{1+f_{D}(x)g^{inter}_{\pm,l}t}.$ (9)
The superconducting critical temperature for each pairing channel is determined from the poles of Eq. 9, which depends on the parity of the C2DEG, as discussed below.
@@
For even values of $J$, the coherence factor $f_{D}(x)=(1-x^{2})/2$ in Eq. 7, arising from the quantum geometric properties of the Haldane phase, leads to the suppression of inter-valley pairing. The critical temperature for the intra-valley (S) pairing (with $f_{S}(x)=1$) is $T^{S}_{c}(x)=\omega_{D}\exp(-1/|g^{intra}_{+,l}|)$, higher than that of the inter-valley (D) pairing, $T^{D}_{c}(x)=\omega_{D}\exp(-2/((1-x^{2})|g^{inter}_{+,l}|))$. As shown in Fig. 4 a), the ratio $T^{D}_{c}/T^{S}_{c}$ is exponentially suppressed, indicating that for even parity the intra-valley pairing dominates over the inter-valley pairing for all values of $x$. This exponential dependence is a direct result of the ratio of the even parity inter- and intra-valley pair susceptibilities $\Pi_{D}/\Pi_{S}$ as derived in Eq. 3.
@@
Since intra-valley pairing dominates for even values of $J$ we focus on the interaction matrix elements associated with $G_{4}({\bf p},{\bf p}^{\prime})$ and $G_{3}({\bf p},{\bf p}^{\prime})$. For even chirality $J$, $(g_{+,l}>g_{-,l})$, with $G_{+,0}=-G_{0}(1-x^{2})$, $G_{+,J}=-G_{0}(1-x)/2$ and $G_{+,-J}=-G_{0}(1+x)/2$. As $G_{+,-J}>G_{+,0}>G_{+,J}$ for $x_{c}>0.5$ we find that the chiral-$J$ Kekulé superconductor has the largest $T^{S}_{c}$ when $k_{F}\leq(\sqrt{3}m/\zeta_{J})^{1/J}$ corresponding to the densities $n_{e}\leq n_{e,2}$ with $n_{e,2}=1/(2\pi)(\sqrt{3}m/\zeta_{J})^{2/J}$. The superconducting order parameter is given by $\Delta(\mbox{\bf r})=\Delta\cos(2{\bf K}_{D}\cdot\mbox{\bf r}+\theta)$ (where $\theta$ corresponds to the relative phase difference between the valleys). Based on the minimal chiral model, we obtain a phase diagram of superconductivity in Fig. 4 b) for even parity chiral 2DEGs in the $(n_{e},m)$ parameter space for different values of $J=4,6$ and $8$. To ensure a simply connected Fermi surface, we assumed that the electron density $n_{e}\geq 0.8\times 10^{12}$ cm-2, sufficiently large [45, 42] to avoid the three-pocket and annulus Fermi surfaces induced by trigonal warping. Due to the large superfluid stiffness of C2DEGs in the dispersive regime [63], we anticipate the Berezinskii–Kosterlitz–Thouless (BKT) transition temperature $T^{S}_{BKT}$ to approach the critical mean field temperature $T^{S}_{BKT}\to T^{S}_{c}$. As detailed in the End Matter, the mean-field critical temperature of the chiral-$J$ Kekulé state is estimated to be in the range $T^{S}_{c}\sim 90.8\textrm{mK}-13.5\textrm{K}$.
@@
Figure 4: a) The ratio of the inter-valley critical temperature ($T_{c}^{D}$) to the intra-valley critical temperature ($T^{S}_{c}$) as a function of $x$. b) Phase diagram for the even chirality C2DEG model of $J$-layer RG with $J=4,6$ and 88 in increasing order. The electron density $n_{e}$ is in units of cm-2 and the Haldane mass is in meV. At low densities, $n_{e}<n_{e,2}$ the chiral-$J$ Kekulé superconducting channel dominates. c) Phase diagram for the odd chirality C2DEG model of $J$-layer RG with $J=3,5$ and 77 in increasing order, with the same axis label as in b). The chiral-$J$ Kekulé superconducting channel dominates for $n_{e}<n_{e,2}$, giving way to ${\bf Q}=0$ valley triplet chiral superconductivity for $n_{e,1}<n_{e}<n_{e,2}$, with the phase boundary indicated by the solid line. Conventional valley singlet s-wave superconductivity dominates for $n_{e}>n_{e,2}$, with the dashed line indicating the phase boundary. d) Schematic of the Kekulé superconductor on the projected honeycomb lattice, with the superconducting unit cell indicated by the dotted lines.
The situation is quite different for odd parity chiral 2DEGs, as the intra-valley pairing only dominates the inter-valley pairing when $x>1/\sqrt{3}$. Therefore, the chiral Kekulé superconducting order has the highest $T_{C}$, when $n_{e}\leq n_{e,1}$ with $n_{e,1}=1/(2\pi)(\sqrt{2}m/\zeta_{J})^{2/J}$. Since $G_{-,-J}=-G_{0}/2(1+x)$ is the dominant channel for odd parity, the intra-valley pairing order exhibits an $\pi$-phase resulting in a Kekulé superconductor with the order parameter $\Delta(\mbox{\bf r})=\Delta\sin(2{\bf K}_{D}\cdot\mbox{\bf r}+\theta)$. $G_{-,J}$ remains the dominant channel for $1/2<x<1/\sqrt{3}$, resulting in a valley-triplet $\pi$-phase chiral-$J$ inter-valley superconducting order for the density regime $n_{e,1}<n_{e}<n_{e,2}$. However, when $n_{e}>n_{e,2}$, $G_{+,0}=-G_{0}(1-x^{2})$ becomes the dominant channel, we get a conventional valley singlet s-wave superconductor with ${\bf Q}=0$. Fig. 4 c) shows the phase diagram for the odd parity chiral 2DEGs as a function of $(n_{e},m)$.
@@
Interestingly, for both parities, the presence of a valley degree of freedom permits even-order pairing channels, including a momentum-independent $s$-wave pairing when $\mu>m$. However, the $s$-wave pairing channel is suppressed as $\mu\to m$ given $G_{+,0}=-G_{0}(1-x^{2})$. This suppression originates from a combination of the Pauli exclusion principle and the quantum geometric properties of the Haldane phase. At the Dirac points $(\pm{\bf K}_{D})$, the sublattice and valley degrees of freedom become coincident and act as a single pseudospin, thereby forbidding local (contact) interactions and rendering $s$-wave Kekulé order incompatible. Consequently, this enhances the relative stability of the chiral-$J$ Kekulé order, which emerges as the dominant channel in this regime. Moreover, due to the presence of a non-zero pair tunnelling channel $G_{3,l}$, we expect the chiral-$J$ Kekulé state to be robust to the application of a displacement field, so long as $\Delta_{V}<m$.